CRY+VIS Interface Meeting
Date: 2021/7/2 13:00-14:10
Participants: Takahashi, Ikeda, Miyo, Yokozawa, Ushiba, Aso, Kimura, Sato, Hirata
- IX: Broken cables of TF/BF actuator were repaired.
- IY: The ballasts on the cryo-payload and the BF were adjusted. Additional spring was implemented on the BF. Height adjustment was scheduled in the next week.
- EX: Alignment for the green beam is planed in the next week.
- EY: Mirror cleaning is on going in Mitaka. 54-deg water was not effective. Absorption measurement will be applied on 7/5. Transport was scheduled on 7/12 at earliest.
- Corrugate tube fixing jig: Implementation is scheduled on 7/12-13.
- TM height setting: Next target is ITMY.
- Commissioning in air: X-arm.
- Commissioning in vacuum:
- Overall commissioning:
- Acceptance check: Group work is set every Friday.
- PF actuator: After SC.
- Jig for FC painting:
- Separation of rack in 2F: This was scheduled on 7/16..
- Disassembly of PRM/PR2 assembly frame: It was carried out on 6/28.
- New vacuum pumps for V-chambers: Check the port location and interference of cabling.
- Extension of cables for picomotor drivers: Test is planed in EYV.
Participants: Takahashi, Fabian, Ishizaki, Tanaka, Sugiyama, Terrence, Aso, Hirata, Sato
This is generated by python script in /users/Miyo/dropbox/Script/scrapping/main.py.
- Adjusted the ballasts on the cryo-payload and the BF in IY. Additional spring was implemented on the BF.
- Repaired the broken cables for TF/BF actuator opening the chamber. The BF and TF were recovered and the chamber was closed.
- We locked IPs and F0 in BS, SR2 and SR3 and Terrence measured LVDT sensitivities.
- Hirata-san measured differewnt parameters for the acceptance check.
- We measured the height of the BS when the GAS Filters are at nominal position (zero): it is 3 mm below the nominal height.
- In case we want to correct this discrepancy it might be good to make reference markings at the +Y and -X sides of the chamber. There's only one marking accesible.
- We should use the jacks to lift it.
BS supporting beams
- It seems no one has a qualitative argument for the installation of the beams, only impressions arising from observations.
- We should ask the people who designed the frame. I guess they put the beams in the design for a reason.
- In the meantime I guess we can just go ahead with commissioning. If later on we realize we need to install the beams, then we do it.
- I applied a sinusoidal actuation at 200 mHz to TM-Y and set the position of the length sensing QPD such that the peak at such a frequency was minimum in the channel measuring horizontal displacement. The klog is pending.
- Then we calibrated the tilt and length sensing QPD readouts in terms of the QPD displacement in the vertical and horizontal directions.
I couldn't make sense of the geophone transfer function for sensor diagonalization and it turned out that the geophones were not probably calibrated. I found out by comparing the geophone response and the LVDT yaw response and found that the geophones have strange responses at ~1 Hz. Turned out the calibration filters were not correct (for H2 and H3). I plugged those from Fujii, which were obtained two years ago. But the H2 geophone response was still bad so I had to manually tweak the calibration filter, which is not the best solution. In the future, we should recalibrate those geophones. See Klog here.
I have settled on a sensing matrix and actuation matrix for the IP. I can't really do much better because the geophones were prefiltered by high pass filters, which distort the phase response of the geophones at lower frequencies (~ IP fundamental resonances), which makes the sensor coupling frequency-dependent. See Klog here
The LVDT and geophone noises were modeled and were used to design complementary filters using my H-infinity method. Klog here. The complementary filter is optimal in the sense that it makes the super sensor noise slightly worse (logarithmic sense) than the lower bound of the sensor noise. The super sensor noise follows the shape of the lower bound in the log-log plot. I have compared my filters with Joris' and Sekiguchi's for the longitudinal DoF and there's roughly two times reduction in the expected RMS. I have installed the filters in the system along with Joris' and Sekiguchi's and then simultaneously measure the free swing spectrum of the IP longitudinal displacement and the result pretty much agrees with my calculation. I can use the same method to design sensor correction filters and damping filters and this is what I will try to do later.
- There are some caveats using my methods. I couldn't close the IP loop using Miyo's damping filters at the beginning. After some debugging, I found that Foton misinterpreted one of my filters and so the complementary filters were not complementary. I have somehow reinstalled the filter manually. It's much better but it's still not entirely complementary. I need more time to investigate what went wrong in the digital system. Anyway, I can close the loop. There's still something strange, but I won't go into the subtle details. (I have a feeling that we might not be able to use the geophones because it's so hard to make it right.)
- Type-A team: Adjustment of IYV
- Type-B team: MSB
IMC suspensions are ready in air  will finish 2021/04/21.
PRs are ready in air  will finish 2021/07/09.
ETMX is ready in air  will finish 2021/08/25.
ITMX is ready in air  will finish 2021/08/31.
ETMY is ready in air  will finish 2021/09/16.
ITMY is ready in air  will finish 2021/11/08.
BS is ready in air  will finish 2021/11/24.
SRs are ready in air  will finish 2021/12/20.
Output suspensions are ready in air  will finish 2021/12/21.
TMSY is ready in air  will finish 2021/??/??.
TMSX is ready in air  will finish 2021/??/??.
The week from 7/5
- Takahashi: 7/5-8 (KIW8 7/7-9)
- Sato: 7/5-8
- Hirata: 7/5-9
- Yano: 7/5-9
The week from 7/12
- Takahashi: 7/12-16
- Sato: 7/12-16
- Hirata: 7/12-16
- Yano: 7/12-16